OpenStudy (anonymous):

What are the solutions of 3x2 + 4x = -5?

OpenStudy (anonymous):

a) -4 +- i sqrt66 over 6 b)-2 +- 2i sqrt66 over 3 c)-2 +- i sqrt11 over 3 d)-4 +- i sqrt11over6

OpenStudy (anonymous):

this is interesting

hartnn (hartnn):

no direct answers, if you want to learn, many are willing to explain.

OpenStudy (anonymous):

ya im willing to learn but if your cool with giving me the answer in the end

OpenStudy (swissgirl):

Ya totalllyyyyy

OpenStudy (swissgirl):

I am all for that :)

OpenStudy (swissgirl):

So lets bring over the 5 to the other side so \(3x^2+4x+5=-5+5\) \(3x^2+4x+5=0\) Now we need to factor

OpenStudy (swissgirl):

Sooooo A=3 B=4 and C=5 Now we need to use the Quadratic Formula Are you Familiar with that?

OpenStudy (swissgirl):